OpenStudy (anonymous):

An ammeter deflects full scale with a current of 5 amperes and has a total resistance of 0.5 ohms what sunt resistance must be connected to it to measure 25 amperes full scale?

OpenStudy (y2o2):

\[Rs = {I_g . R_g \over I-I_g} = {5 \times 0.5 \over 25-5} = {1 \over 8} \ \ ohm \]

OpenStudy (radar):

Lets think about this. If we are attempting to measure 25 amps, and the ammeter is full scale at 5 amps, then the shunt must accomodate the excess 20 amps. Now the voltage drop this shunt must equal the drop as when 5 amps goes through the .5 ohm meter resistance which would be .5 X 5 or 2.5 volt. What resistance would produce 2.5 volts when 20 amps is flowing through it? E=IR 2.5=20R R=2.5/20=.125 ohm or 1.8 as y2o2 posted above.